BerkeleyNetworkQuickReference May1979 (updatedMarch1980) CommandSummaryExampleExplanation netcpfromfiletofile%netcpdefs.hC:defs.hcopydefs.htoCmachine %netcpA:fig1fig1copyfromAtohere %netcpCory:q1D:q1copyfromCorytoD mailmach:user%mailA:alansendmailto‘‘alan’’onA %mailA:alanC:georgemultiplerecipients netmail[−c][mach]%netmailCoryreadmailonCorymachine %netmail−cCory:userncheckmailonCorymachine netlpr[−mmach]file%netlpr−mDp1.pprintp1.ponDlineprinter netq[−a][mach]%netqlistnetworkqueue %netq−aCory...allfilestoCory netrm[−][code]%netrm−removetheuser’srequests net[−mmach][−]command%net−mAwhosendwhotoAmachine %net−−mAlpr<f1lprfile‘‘f1’’onA Thenetcp,netmail,netlpr,andnetcommandsneedremoteloginnamesandpasswords.They maybeprovidedwhenprompted-forontheterminal,onthecommandline,orina‘‘.netrc’’filein theuser’slogindirectory. OtherOptions −lloginloginnameonremotemachine −ppasswordpasswordonremotemachine −fpromptforloginnameandpassword −nmailbackratherthanwriteback −qsilentmode−nomessagessentback −mmachsendtomach(onlyfornet,netlpr) −inputfromstandardinput(onlyfornet) Example.netrcfile: defaultZ machineYlogindracula machineZloginschmidt,quietyes .netrcoptionsforeachmachine OptionParameterDefaultComment loginnamelocalnameloginnameforremotemachine passwordpassword(none)passwordforremoteloginname commandcommand(none)defaultcommandtobeexecuted writeyes/noyesifpossible,writetouser forceyes/nonoalwayspromptfornameandpassword quietyes/nonolikethe−qoption BerkeleyNetworkQuickReference,Page2 TheCurrentStateoftheBerkeleyUNIXNetwork MachineInternalRunDefaultOther NameNameByMachineName(s) AAComputerCenterC BBComputerCenterD CCComputerCenterA DDComputerCenterC EEComputerCenterC Ing70IINGRESGroupIngVAXIngres IngVAXJINGRESGroupIng70 ImageMSakrisonESVAX ESVAXOEE-CEResearchCSVAX SRCSSurveyRes.Cent.D CSVAXVCSResearchCory CoryYEECSDept.CSVAX EECS40ZEECSDept.ESVAX Topology ThelinksbetweenIng70−IngVAX,Ing70−CSVAX,A−C,C−D,C−E,andB−Drunat9600 Baud,theotherlinksrunat1200Baud. FreeCommands(loginas‘‘network’’,nopassword). bpq netlogrcsvpqwhom epq netqrcslogwwrite finger psrcsqwhereyank lpq pstattrqwho Mailisfreetoanymachine. netlpr isfreebetweenComputerCentermachines. The lpr commandisfreeontheINGRESmachines. TheEECS40hasnofreecommands(exceptforremotemail).